The length of a rectangle is 8 in more than twice its width. If the perimeter of the rectangle is 34 in, find the width of the rectangle.
Answer by checkley77(12844)

L=2W+8
2(2W+8)+2W=34
4W+16+2W=34
6W=34-16
6W=18
W=18/6
W=3 FOR THE WIDTH.
L=2*3+8=6+8=14 FOF THE LENGTH.
PROOF:
2*3+2*14=34
6+28=34
34=34